AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 458     Entry time: Fri Jan 5 12:29:10 2024
Author: JM, CB 
Subject: Further bias tests 

Biased detectors this morning to 150V, the leakage current for all detectors appears stable and is ~3.5 uA. We have also replaced the CAEN bias module so that all channels are now functional. Detectors are connected as follows;

CHN0 -> Top left

CHN1 -> Top Right

CHN2 -> Bottom Left

CHN3 -> Bottom Right

Reconnected the motors and moved the detectors while biased. No observable change in the leakage current during movement at 150V. An increase in the pressure of the electron target was observed during movement (up to 1e-9 mbar). This increase was only momentarily, and the pressure recovered after a couple of seconds. We expect the pressure increase from movement to decrease over time.

 

ELOG V3.1.4-unknown